OpenStudy (anonymous):

solve: 3(8n-5)^1/3-9 = 0

OpenStudy (anonymous):

\[3(8n-5)^{1/3} -9=0\]

OpenStudy (souvik):

add 9 in both sides.. divide by 3 now cube both sudes

OpenStudy (nurali):

\[3(8n-5)^{1/3}=9\] \[(8n-5)^{1/3}=3\] Multiply Power 3 on both sides \[8n-5=3^{3}\] \[8n-5=27\] \[8n=32\] \[n=4\]

OpenStudy (anonymous):

thank you !

OpenStudy (nurali):

My pleasure.